5a+3= 31-2a can you help?
Found 2 solutions by ikdeep, sdmmadam@yahoo.com:
Answer by ikdeep(226)

5a+3= 31-2a
5a + 2a = 31 - 3
7a = 28
or a = 28/7
a = 4 .............answer

Answer by sdmmadam@yahoo.com(530)

5a+3= 31-2a ----(1)
That is 5a+2a = 31-3 (adding 2a to both the sides and subtracting 3 from both the sides, or grouping like terms changing sign while changing side)
7a = 28
a = 28/7 = 4
Answer: a = 4
Verification: Putting a = 4 in (1)
LHS = 5a + 3 = 20+3 = 23
RHS = 31-2a = 31-8 = 23 = LHS
Therefore our value is correct.